Calculate the consumer surplus for the demand function P=48-0.2Q when the market price is P=30
Answer by checkley79(3341)

P=48-0.2Q
30=48-0.2Q
0.2Q=48-30
0.2Q=18
Q=18/0.2
Q=90 ANS.